# Finite and full scale localization for the multi-frequency quasi-periodic CMV matrices

### v2

@@ 6 Full Scale Localization
Let $\varepsilon\in(0,\frac{1}{5})$ and $\hat{\Omega}_{N_{0}}$, ^$\hat{\mathcal{B}}_{N_{0},\omega}$ be as in Lemma 9.6 For any $N_{0}\geq C(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\hat{\Omega}_{N_{0}}$, $x\in\mathbb{T}^{d}\backslash\hat{\mathcal{B}}_{N_{0},\omega}$, the following statements hold. Let $n_{k}=n^{2^{k}}$. If $L(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x))>2\gamma$ and $I=I(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x),x)\subset[-\frac{n-1}{2},\frac{n-1}{2}]$, then for each $k\geq 1$ there exist $j_{k}$ such that
@@
From (6.9), we know that (6.6) holds for $n_{k-1}-1\leq|s|\leq n_{k}-1$. Then it is enough to prove that (6.6) holds for $\frac{3(n-1)}{4}\leq|s|\leq n_{k-1}$. For $21\leq\hat{k}\leq k-$ 2, due to
@@
for $n_{\hat{k}-1}-1\leq|s|\leq n_{\hat{k}}-1$. Due to $n_{k}=n^{2^{k}}$, $21\leq\hat{k}\leq k-$ 2, we have $n-1\leq n_{\hat{k}-1}-1\leq|s|\leq n_{\hat{k}}-1\leq n_{k}-1$. Thus, (6.6) holds for $n-1\leq|s|\leq n_{k}-1$. According to Lemma 6.4,
@@ 7.3 Subharmonic Functions
[13, Lemma 2.4] Let $u(x_{1},\ldots,x_{d})$ be a separately subharmonic function on $\mathbb{T}_{h}^{d}$. Suppose that $\sup\limits_{\mathbb{T}_{h}^{d}}u\leq M$. There are constants $C_{d}$, $C_{d,h}$ such that, if for some $10<\delta<$ 1 and some $L$
@@ 7.4 Cartan’s Estimate
If $d\geq 2$ is an integer and $\mathcal{B}\subset\mathop{\prod}\limits_{j=1}^{d}\mathcal{D}(z_{j,0},1)\subset\mathbb{C}^{d}$, then we define inductively that $\mathcal{B}\in\mathrm{Car}_{d}(H,K)$ if for any $1\leq j\leq d$ there exists $\mathcal{B}_{j}\subset\mathcal{D}(z_{j,0},1)\subset\mathbb{C}$, $\mathcal{B}_{j}\in\mathrm{Car}_{1}(H,K)$ so that $\mathcal{B}^{(j)}_{z}\in\mathrm{Car}_{d-1}(H,K)$ for any $z\in\mathbb{C}\backslash\mathcal{B}_{j}$, here $\mathcal{B}^{(j)}_{z}=\{(z_{1},\ldots,z_{d})\in\mathcal{B}:z_{j}=z\}$.
@@
for any $z\in\frac{1}{6}\mathcal{P}\backslash\mathcal{B}$. Furthermore, when $d=1$ we can take $K=C(M-m)$ and keep only the disks of $\mathcal{B}$ containing a zero of $\varphi$ in them.
$\mathbf{Acknowledgments}$
This work was supported in part by the NSFC (No. 11571327, 11971059).

### v3

@@ 6 Full Scale Localization
Finally, we establish full scale localization by iterating the eigenpair convergence (Lemma 6.6) over a sequence of exponentially expanding intervals $n_{k}=n^{2^{k}}$. We show that the eigenpair sequence converges to an eigenfunction of the infinite MF-QP CMV matrix that decays exponentially everywhere, confirming pure point spectrum and completing the localization proof.
@@
Let $\varepsilon\in(0,\frac{1}{5})$ and $\hat{\Omega}_{N_{0}}$, $\hat{\mathcal{B}}_{N_{0},\omega}$ be as in Lemma 6.6. For any $N_{0}\geq C(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\hat{\Omega}_{N_{0}}$, $x\in\mathbb{T}^{d}\backslash\hat{\mathcal{B}}_{N_{0},\omega}$, the following statements hold. Let $n_{k}=n^{2^{k}}$. If $L(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x))>2\gamma$ and $I=I(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x),x)\subset[-\frac{n-1}{2},\frac{n-1}{2}]$, then for each $k\geq 1$ there exist $j_{k}$ such that
@@
From (6.9), we know that (6.6) holds for $n_{k-1}-1\leq|s|\leq n_{k}-1$. Then it is enough to prove that (6.6) holds for $\frac{3(n-1)}{4}\leq|s|\leq n_{k-1}$. For $1\leq\hat{k}\leq k-2$, due to
@@
for $n_{\hat{k}-1}-1\leq|s|\leq n_{\hat{k}}-1$. Due to $n_{k}=n^{2^{k}}$, $1\leq\hat{k}\leq k-2$, we have $n-1\leq n_{\hat{k}-1}-1\leq|s|\leq n_{\hat{k}}-1\leq n_{k}-1$. Thus, (6.6) holds for $n-1\leq|s|\leq n_{k}-1$. According to Lemma 6.4,
@@ 7.3 Subharmonic Functions
[13, Lemma 2.4] Let $u(x_{1},\ldots,x_{d})$ be a separately subharmonic function on $\mathbb{T}_{h}^{d}$. Suppose that $\sup\limits_{\mathbb{T}_{h}^{d}}u\leq M$. There are constants $C_{d}$, $C_{d,h}$ such that, if for some $0<\delta<1$ and some $L$
@@ 7.4 Cartan’s Estimate
[…]
@@
[…]
Acknowledgments
D.P was supported in part by the NSFC (No. 11571327, 11971059). B.Z was supported by Nankai Zhide Foundation.